OpenStudy (anonymous):

Translate the following sentence into an equation. Do not solve. “Three more than one half a number is 21.”

OpenStudy (anonymous):

(x/2) + 3 = 21

OpenStudy (phi):

Three 3 More than. + One half. 1/2 A number. X Is. = 21 3+ (1/2)x= 21

OpenStudy (anonymous):

thanks